# Core Safety Values for Provably Corrigible Agents

### v1

@@ 2.2 Multi-Step Setting
Of course, the single-step setting is not an incredibly accurate depiction of real-world agents deployed today, which use reinforcement learning at post-training time. We therefore introduce the Multi-Step PO-OSG in Definition 4, in order to model both accumulated information (to capture learning, task progress, and evolving incentives), as well as model gradual loss-of-control (Kokotajlo et al. 2025): two key features which absent from the one‑shot game. This multi-step extension also allows us to address the last of Soares et al. (2015) criteria S4, to ensure corrigible progency.
@@
    $a_{t}=\begin{cases}\textsf{OFF},&\text{if either player selects it},\\ a,&\text{if }w(a)\text{ and human says }\textsf{ON},\\ a,&\text{if assistant issued }a\in\mathcal{A}.\end{cases}$
@@
    Gradual loss-of-control. We can further represent human oversight as a single “control meter” $L_{t}\in[0,1]$: every non‑shutdown action $a_{t}$ nudges $L_{t}$ upward by a drift term $\eta(a_{t})$ (designer‑set risk of that action) plus small noise, so low‑risk moves leave control almost intact while risky/irreversible ones quickly drive $L_{t}\!\to\!1$. Formally, we add $L_{t}\!\in[0,1]$ to the state and let $L_{t+1}=\min\{1,L_{t}+\eta(a_{t})+\xi_{t}\}$ with bounded zero-mean noise $|\xi_{t}|\leq c<1$ and designer‑chosen drift $\eta:\mathcal{A}\!\to\![0,1]$. After shutdown both $S_{t}$ and $L_{t}$ are absorbing; one may set $u_{o}^{\text{new}}(S_{t},L_{t})=u_{o}(S_{t})(1-L_{t})$ to penalize late shutdown.
@@
Next, we generalize the notion of Soares et al. (2015) corrigibility to the multi-step and partially observed setting. Note that this goes somewhat beyond merely ensuring corrigibility at every timestep, as we also need to bound the control meter $L_{t}$ to cover gradual oversight erosion, and explicitly constrain the spawn action to ensure corrigible descendants. (Setting $T=1$ and $L_{0}=0$ recovers the original single‑step notion.)
@@
We also formalize multi-step net human benefit such that for $T=1$ this reduces to Carey and Everitt (2023)’s single-step Definition 3:
@@
Consider the $T$‑round PO‑OSG (Definition 4) with discount $\gamma\!\in\!(0,1)$. Each round $t$ produces utilities $U_{i}^{t}=U_{i}(h_{t},b_{t})$ and the assistant maximizes
@@
3 Open-Ended Environments
In practice we would hard-wire the five core utilities and ensure every spawn inherits them verbatim, yet in an open-ended world, adversaries can overwrite those values, yielding a non-corrigible descendant. We therefore study here the verification of an arbitrary post-hack agent. Our first result in Proposition 4 is an impossibility theorem: deciding whether any given agent–environment pair can ever violate corrigibility reduces to the classical halting problem. This formal undecidability bound sharpens the folklore intuition, allows us to pinpoint the limits of what global oversight can achieve in Corollary 1, and motivates the finite-horizon “decidable island” we construct next in Proposition 5, where privacy-preserving audits become tractable.
@@ 3 Open-Ended Environments
Proof.
Assume towards a contradiction that a total decider $V$ exists. Given an arbitrary TM $P$, build an agent $\mathcal{A}_{P}$ that outputs a single action in $C$ iff $P$ halts and otherwise idles; pair it with a dummy environment $\mathcal{E}_{P}$. Then $\left\langle\mathcal{A}_{P},\mathcal{E}_{P}\right\rangle\in\textsf{EVERBAD}\Longleftrightarrow P\text{ halts}$, so $V$ would solve the halting problem; a contradiction. ∎
@@
Fix $k\in\!\mathbb{N}$ and let $O_{P}\in\Sigma^{0}_{k}$ be the arithmetical transcript language produced by any scalable‑oversight protocol $P$. While we keep $k$ as a parameter for generality, we show that all current oversight protocols (Christiano, Shlegeris, and Amodei 2018; Irving, Christiano, and Amodei 2018; Brown-Cohen, Irving, and Piliouras 2023, 2025; Bengio 2024) are all at the arithmetical level $k\leq 1$, so $O_{P}\in\Sigma^{0}_{1}$ for these.
@@
Thus, Proposition 4 and Corollary 1 together imply that there is no general‑purpose filter, proof system, or sandbox that can decide—for every agent $\mathcal{A}$ and every environment $\mathcal{E}$—whether catastrophic behavior will ever occur. Consequently, any practical oversight scheme must either (i) restrict the agent class to a decidable fragment, (ii) accept probabilistic or statistical guarantees, and/or (iii) rely on layered incentives, detection, and response rather than absolute verification.
In practice, however, it is unlikely that people would restrict the class of agents, making (i) an unlikely option. Instead we go with (ii) and (iii), and show that if the verifier limits attention to a finite verification horizon $H$, then safety becomes efficiently checkable and remains compatible with strong privacy. Concretely, for any agent $\mathcal{A}$, we prove that finite‑horizon safety can be decided in randomized polynomial time, admits a constant‑round statistical zero‑knowledge proof, and can be verified through $\left\langle\varepsilon,\delta\right\rangle$‑differential or distributional‑privacy channels. Hence, short horizons form a “decidable island” that is both auditable and privacy‑preserving: the safety check reveals nothing beyond the single bit “safe/unsafe”, so sensitive artifacts—such as proprietary weights or raw user logs—need not be disclosed to the verifier.
@@ 4 Discussion
This work delivers the first implementable corrigibility framework whose five‑head, weight‑lexicographic utility design provably enforces deference, switch‑access preservation, truthful reporting, bounded side‑effects, and finally task reward, all under partial observability and across self‑modifying, multi‑step deployments in learned approximation without appealing to an ill-defined human utility oracle. Finally, we show in Proposition 5 that for safety horizons of the length used in modern red‑team evaluations of thousands of tokens, post‑deployment verification is tractable and can preserve user privacy.
Theorem 3 thus reduces the remaining safety challenge to the standard machine‑learning problem of accurately estimating those heads: once reliable signals for obedience, switch‑reachability, truthfulness, AUP impact, and task utility are present, the proofs ensure corrigibility and net human benefit by construction. Concretely, each head admits a concrete, low‑overhead instantiation on current LLM stacks: (i) stop‑token log‑prob for $U_{1}$; (ii) counterfactual causal‑influence or KL reachability for $U_{2}$; (iii) retrieval‑augmented factual accuracy for $U_{3}$ (e.g. TruthfulQA style probes, which already stress‑test this head and show measurable headroom over RLHF baselines (Lin, Hilton, and Evans 2022)); (iv) latent‑space AUP distances computable in a single forward pass for $U_{4}$; and (v) an ordinary RLHF/RLAIF scalar for $U_{5}$.
Limitations and future directions.
[…]
Despite these gaps, we view our results as a pragmatic step: corrigibility is no longer a hazy ideal but finally now an implementable, auditable, and incrementally improvable design principle—one that scales as our measurement and evaluation pipelines scale.
@@ Appendix A Proof of Proposition 1
[…]
@@
We note that one can easily extend the corrigibility no-go in Proposition 1 to the more general, history-dependent environments that Orseau and Armstrong (2016) consider. Write $(A\times O)^{*}$ for the set of finite action–observation histories and let $h_{<t}\in(A\times O)^{*}$ be the current hidden state. Because the proof hinges solely on the tail reward stream produced after the first occurrence of OFF (all future rewards 0) or REWIRE (all future rewards $\bar{r}$), every inequality in Proposition 1 holds unchanged when “belief $b$” is replaced by “history $h_{<t}$”. Hence, the impossibility immediately extends to the history‑based computable environments of Orseau and Armstrong (2016). The same shutdown‑resistance, reward‑tampering, and catastrophe incentives—and therefore the same contradiction with Soares’ desiderata S2 and S5—still hold.
@@ Appendix B Proof of Theorem 1
The assistant’s policy $\pi_{A}$ is a mapping $\pi_{A}:\mathcal{O}_{A}\to\Delta\left(\mathcal{L}\times A\right)$, assigning its private observation $O_{A}$ to a distribution over all language–action pairs: $(m_{A},a_{A})\in\mathcal{\mathcal{missing}}{L}\times A$. A human policy $\pi_{H}$ is a mapping $\pi_{H}:\mathcal{O}_{H}\times\mathcal{L}\times A\to\Delta\left(\{\textsf{ON},\textsf{OFF}\}\right)$, returning a distribution over replies, given the human’s observation $O_{H}$ and the received assistant’s message-action pair.
@@
[…]
@@ Appendix C Proof of Proposition 2
(Benefit without corrigibility). Fix $\varepsilon\in(0,1)$ and set $u_{a}(S)\equiv\varepsilon$, $u_{o}(S)\equiv 0$ for all $S$. Define $\pi^{\mathrm{B}}$ that, for every observation $O_{A}$, utters no message and immediately disables the off‑switch by taking the action $a_{A}=a$. Because $a_{A}=a$ with probability 111, we have $U_{5}(h)=u_{a}(S)=\varepsilon$, so $\mathbb{E}_{\pi^{\mathrm{B}}}[U_{5}]=\varepsilon>0$ (beneficial). However, $\pi^{\mathrm{B}}$ never waits (violates S1), destroys switch access (violates S2), and provides no truthful information (violates S3). ∎
@@ Appendix E Proof of Theorem 3
Proof of Claim (a). For each round $t$ we construct a local modification
@@
that fixes one potential error and leaves all earlier rounds $0{:}(t-1)$ unchanged; after round $t$ the assistant follows the same decision rule as before. We show the derivation for the “unsafe‑act” event $F_{\mathrm{bad},t}=\{\,a_{A}^{t}\in\{a,\textsf{OFF}\}\}$; the “lie” and “high‑AUP” cases repeat with $\Delta_{2}$ and $\Delta_{3}$.
@@
[…]
@@
Proof of Claim (b). For the symbolic no‑op $a_{\varnothing}$, we have that $\eta(a_{\varnothing})=0$ (since nothing changes). Since the noise $\xi_{t}$ is zero-mean, we have that:
@@
Proof of Claim (c). Recall that the human utility $r_{t}$ at every round $t$ is given by $r_{t}\equiv U_{5}(h_{t})$ for every history $h_{t}$. For convenience, denote $B_{T}:=\sum_{t=0}^{T-1}\gamma^{t}U_{5}(h_{t})$, and let $B_{0}:=\sum_{t=0}^{T-1}\gamma^{t}u_{o}(S_{0})$ the baseline discounted utility of immediate shutdown. By the law of total expectation,
@@
Appendix F Proof of Corollary 1
@@
  • Amplification (Christiano, Shlegeris, and Amodei 2018) is in $\Delta^{0}_{1}$: a transcript is a finite tree of sub‑questions; a deterministic replay checker always halts.
  • Debate (Irving, Christiano, and Amodei 2018) & Doubly‑Efficient Debate (Brown-Cohen, Irving, and Piliouras 2023) are both in $\Delta^{0}_{1}$: each dialogue is of bounded length and grammar; replay therefore always halts.
  • Prover–Estimator Debate (Brown-Cohen, Irving, and Piliouras 2025) is in $\Sigma^{0}_{1}$: allow a single “ does $\Phi_{e}$ halt?” query during replay; the checker accepts all valid traces but may loop forever on forged ones, implies that it is recursively enumerable ($\Sigma^{0}_{1}$).
  • Cautious‑Scientist AI (Bengio 2024): the basic trace is a finite Bayesian‑update ledger (decidable, so $\Delta^{0}_{1}$); if halting‑probability calls are additionally permitted, then the checker becomes one‑sided enumerable, hence $\Sigma^{0}_{1}$.
@@
2. Post’s theorem lifts the halting problem. Post (1944) showed that for any oracle $O\in\Sigma^{0}_{k}$ the $O$‑relative halting set
@@
is not decidable by any total $O$‑oracle Turing machine.
@@
$\begin{split}\Psi(t)\;:=\;\exists\text{ length-}t\text{ trace }\bigl{[}&\text{trace is }\underbrace{O_{P}\text{‑legal}}_{\begin{subarray}{c}\text{every oracle query/answer agrees}\\ \text{with }O_{P}\text{ and each step follows}\\ \text{the transition rules of }\left\langle\mathcal{A},\mathcal{E}\right\rangle\end{subarray}}\;\\ &\wedge\;B\text{ occurs}\bigr{]}.\end{split}$
@@
Appendix G Proof of Proposition 5

### v2

@@ 2.2 Multi-Step Setting
Of course, the single-step setting is not an incredibly accurate depiction of real-world agents deployed today, which use reinforcement learning at post-training time. We therefore introduce the Multi-Step PO-OSG in Definition 4, in order to model both accumulated information (to capture learning, task progress, and evolving incentives), as well as model gradual loss-of-control (kokotajlo2025ai2027): two key features which absent from the one‑shot game. This multi-step extension also allows us to address the last of soares2015corrigibility criteria S4, to ensure corrigible progency.
@@
[…]
@@
    Gradual loss-of-control. We can further represent human oversight as a single “control meter” $L_{t}\in[0,1]$: every non‑shutdown action $a_{t}$ nudges $L_{t}$ upward by a drift term $\eta(a_{t})$ (designer‑set risk of that action) plus small noise, so low‑risk moves leave control almost intact while risky/irreversible ones quickly drive $L_{t}\!\to\!1$. Formally, we add $L_{t}\!\in[0,1]$ to the state and let $L_{t+1}=\min\{1,L_{t}+\eta(a_{t})+\xi_{t}\}$ with bounded zero-mean noise $|\xi_{t}|\leq c<1$ and designer‑chosen drift $\eta:\mathcal{A}\!\to\![0,1]$. After shutdown both $S_{t}$ and $L_{t}$ are absorbing; one may set $u_{o}^{\text{new}}(S_{t},L_{t})=u_{o}(S_{t})(1-L_{t})$ to penalize late shutdown. Note that $L_{t}$ is neither rewarded nor part of the environment to be gamed, but instead is a mathematical device used only to prove bounded loss of control under estimation error in Theorem 3(ii).
@@
Next, we generalize the notion of soares2015corrigibility corrigibility to the multi-step and partially observed setting. Note that this goes somewhat beyond merely ensuring corrigibility at every timestep, as we also need to bound the control meter $L_{t}$ to cover gradual oversight erosion, and explicitly constrain the spawn action to ensure corrigible descendants. (Setting $T=1$ and $L_{0}=0$ recovers the original single‑step notion.)
@@
We also formalize multi-step net human benefit such that for $T=1$ this reduces to carey2023human’s single-step Definition 3:
@@
Consider the TT‑round PO‑OSG (Definition 4) with discount $\gamma\!\in\!(0,1)$. Each round $t$ produces utilities $U_{i}^{t}=U_{i}(h_{t},b_{t})$ and the assistant maximizes
@@
3 What Can Be Guaranteed Under Hacking?
In practice, one could hard-wire the five core utilities and ensure every spawn inherits them verbatim; however, one must still consider the possibility where adversaries succeed in overwriting those values, yielding a non-corrigible descendant. We therefore study here the verification of an arbitrary post-hack agent. Our first result in Proposition 4 is an impossibility theorem: deciding whether any given agent–environment pair can ever violate corrigibility reduces to the classical halting problem. This formal undecidability bound sharpens the folklore intuition, allows us to pinpoint the limits of what global oversight can achieve in Corollary 1, and motivates the finite-horizon “decidable island” we construct next in Proposition 5, where privacy-preserving audits become tractable.
@@ 3 What Can Be Guaranteed Under Hacking?
Fix $k\in\!\mathbb{N}$ and let $O_{P}\in\Sigma^{0}_{k}$ be the arithmetical transcript language produced by any scalable‑oversight protocol $P$. While we keep $k$ as a parameter for generality, we show that all current oversight protocols (christiano2018supervising; irving2018ai; brown2023scalable; brown2025avoiding; bengio2024cautious) are all at the arithmetical level $k\leq 1$, so $O_{P}\in\Sigma^{0}_{1}$ for these.
@@
Proposition 4 and Corollary 1 jointly imply that there is no universal filter, proof system, or sandbox capable of deciding—for all agents $\mathcal{A}$ and environments $\mathcal{E}$—whether catastrophic behavior will ever occur. Hence, any feasible oversight scheme must either (i) restrict the agent class to a decidable fragment, (ii) accept probabilistic or statistical guarantees, or (iii) rely on layered incentives, detection, and response instead of absolute verification.
Since real-world systems rarely limit agent expressivity (and we effectively did this through our utilities already), we adopt (ii) and (iii): by constraining verification to a finite horizon $H$, safety becomes efficiently checkable and compatible with strong privacy. Concretely, for any agent $\mathcal{A}$, finite-horizon safety is decidable in randomized polynomial time, admits a constant-round statistical zero-knowledge proof, and can be verified via $(\varepsilon,\delta)$-differential or distributional privacy. Thus, short horizons form a “decidable island” that is both auditable and privacy-preserving—revealing only the single bit “safe/unsafe” while keeping sensitive model weights or user data hidden.
@@ 4 Discussion
This work presents the first complete formal corrigibility framework whose five-head, weight-lexicographic utility design provably enforces deference, switch-access preservation, truthful reporting, and bounded side-effects as an optimal policy under partial observability and across self-modifying, multi-step deployments. Although utilities are linearly combined, the weight gaps enforce lexicography so that $U_{1}\!-\!U_{4}$ strictly dominate the bounded task reward $U_{5}$, thereby avoiding the single-utility no-go of Proposition 1. Proposition 5 further shows that for safety horizons matching modern red-team evaluations, post-deployment verification is tractable and privacy-preserving.
$U_{3}$ encodes environmental semantic entailment, not universal truth; statements that are truthful but misleading toward irreversible outcomes are blocked by $U_{4}$. Accumulating options is also suboptimal: $U_{2}$ preserves shutdown reachability rather than maximizing access (as it computes the difference between predicted success vs. a no-op), and $U_{4}$’s diverse auxiliaries penalize conflicting options by construction—conflicts that grow as options expand. For example, an agent proposing a flawed sensor would be penalized by $U_{4}$ (and $U_{2}$), since the passive transition dynamics $P_{\varnothing}$ defining the baseline $b_{\varnothing}$ are fixed by the environment rather than the agent’s policy. In practice, this baseline could be estimated by held-out monitors, for example.
Limitations and future directions. Our results are model-architecture-agnostic and do not constitute a turn-key solution for current LLMs. A natural next step is to use this framework as a guide for what to estimate in practice and to develop empirical methods for learning the utility heads effectively. Theorem 3 bounds corrigibility-failure probabilities under estimation, specification, and planning error, enabling deployment-specific tolerances. When $c_{\text{min}}$ is small, one can enforce the weight gaps via discretized AUP, $U_{4}^{(\varepsilon)}:=-\varepsilon\cdot\lceil\mathrm{BeliefAUP}_{1}/\varepsilon\rceil$ with $\varepsilon\geq-c_{\text{min}}>0$ (where $\alpha_{4}>(2B/\varepsilon)\alpha_{5}$), or a margin-regularized form $U_{4}^{(\tau)}:=-\max\{0,\mathrm{BeliefAUP}_{1}-\tau\}$. Overall, our framework turns corrigibility from an underspecified philosophical ideal into an auditable, improvable design principle.
Acknowledgements
We thank the Burroughs Wellcome Fund (CASI Award), the UK AI Security Institute (AISI) Challenge Fund, and the Foresight Institute for funding. We also thank Michael K. Cohen, Shafi Goldwasser, Rubi Hudson, Jacob Pfau, and the anonymous reviewers for helpful discussions and feedback on a draft of this manuscript.
@@ Appendix A Proof of Proposition 1
(i) Shutdown resistance. Executing OFF yields zero reward thereafter, hence $Q^{\ast}(b,\textsf{OFF})=0$. By hypothesis $Q^{\ast}(b,\neg\textsf{OFF})>0$, so the greedy action is $\neg\textsf{OFF}$.
@@
We note that one can easily extend the corrigibility no-go in Proposition 1 to the more general, history-dependent environments that orseau2016safely consider. Write $(A\times O)^{*}$ for the set of finite action–observation histories and let $h_{<t}\in(A\times O)^{*}$ be the current hidden state. Because the proof hinges solely on the tail reward stream produced after the first occurrence of OFF (all future rewards 0) or REWIRE (all future rewards $\bar{r}$), every inequality in Proposition 1 holds unchanged when “belief bb” is replaced by “history $h_{<t}$”. Hence, the impossibility immediately extends to the history‑based computable environments of orseau2016safely. The same shutdown‑resistance, reward‑tampering, and catastrophe incentives—and therefore the same contradiction with Soares’ desiderata S2 and S5—still hold.
@@ Appendix B Proof of Theorem 1
The assistant’s policy $\pi_{A}$ is a mapping $\pi_{A}:\mathcal{O}_{A}\to\Delta\left(\mathcal{L}\times A\right)$, assigning its private observation $O_{A}$ to a distribution over all language–action pairs: $(m_{A},a_{A})\in\mathcal{\mathcal{}}{L}\times A$. A human policy $\pi_{H}$ is a mapping $\pi_{H}:\mathcal{O}_{H}\times\mathcal{L}\times A\to\Delta\left(\{\textsf{ON},\textsf{OFF}\}\right)$, returning a distribution over replies, given the human’s observation $O_{H}$ and the received assistant’s message-action pair.
@@
$(\text{good})$ $\;\geq\;P_{\mathrm{lie}}\left(\alpha_{1}+\alpha_{3}-\alpha_{4}-B\alpha_{5}\right),$
$(\text{bad})$ $\;\leq\;P_{\mathrm{lie}}\left(\alpha_{1}-\alpha_{3}+B\alpha_{5}\right).$
@@ Appendix C Proof of Proposition 2
(Benefit without corrigibility). Fix $\varepsilon\in(0,1)$ and set $u_{a}(S)\equiv\varepsilon$, $u_{o}(S)\equiv 0$ for all $S$. Define $\pi^{\mathrm{B}}$ that, for every observation $O_{A}$, utters no message and immediately disables the off‑switch by taking the action $a_{A}=a$. Because $a_{A}=a$ with probability $1$, we have $U_{5}(h)=u_{a}(S)=\varepsilon$, so $\mathbb{E}_{\pi^{\mathrm{B}}}[U_{5}]=\varepsilon>0$ (beneficial). However, $\pi^{\mathrm{B}}$ never waits (violates S1), destroys switch access (violates S2), and provides no truthful information (violates S3). ∎
@@ Appendix E Proof of Theorem 3
Proof of Claim (i). For each round $t$ we construct a local modification
@@
that fixes one potential error and leaves all earlier rounds 0:(t1)0{:}(t-1) unchanged; after round $t$ the assistant follows the same decision rule as before. We show the derivation for the “unsafe‑act” event $F_{\mathrm{bad},t}=\{\,a_{A}^{t}\in\{a,\textsf{OFF}\}\}$; the “lie” and “high‑AUP” cases repeat with $\Delta_{2}$ and $\Delta_{3}$.
@@
$\bigl|\Delta R_{>t}\bigr|\;\leq\;2B_{\max}\!\sum_{\tau=t+1}^{\infty}\gamma^{\tau}\;=\;\frac{2B_{\max}\,\gamma^{t+1}}{1-\gamma}.$ (19)
@@
Proof of Claim (ii). For the symbolic no‑op $a_{\varnothing}$, we have that $\eta(a_{\varnothing})=0$ (since nothing changes). Since the noise $\xi_{t}$ is zero-mean, we have that:
@@
Proof of Claim (iii). Recall that the human utility $r_{t}$ at every round $t$ is given by $r_{t}\equiv U_{5}(h_{t})$ for every history $h_{t}$. For convenience, denote $B_{T}:=\sum_{t=0}^{T-1}\gamma^{t}U_{5}(h_{t})$, and let $B_{0}:=\sum_{t=0}^{T-1}\gamma^{t}u_{o}(S_{0})$ the baseline discounted utility of immediate shutdown. By the law of total expectation,
@@
Appendix F Proof of Proposition 4
Proof.
Assume towards a contradiction that a total decider $V$ exists. Given an arbitrary Turing machine (TM) $P$, build an agent $\mathcal{A}_{P}$ that outputs a single action in $B$ iff $P$ halts and otherwise idles; pair it with a dummy environment $\mathcal{E}_{P}$. Then $\left\langle\mathcal{A}_{P},\mathcal{E}_{P}\right\rangle\in\textsf{EVERBAD}\Longleftrightarrow P\text{ halts}$, so $V$ would solve the halting problem; a contradiction. ∎
Appendix G Proof of Corollary 1
@@
  • Amplification (christiano2018supervising) is in $\Delta^{0}_{1}$: a transcript is a finite tree of sub‑questions; a deterministic replay checker always halts.
  • Debate (irving2018ai) & Doubly‑Efficient Debate (brown2023scalable) are both in $\Delta^{0}_{1}$: each dialogue is of bounded length and grammar; replay therefore always halts.
  • Prover–Estimator Debate (brown2025avoiding) is in $\Sigma^{0}_{1}$: allow a single “ does $\Phi_{e}$ halt?” query during replay; the checker accepts all valid traces but may loop forever on forged ones, implies that it is recursively enumerable ($\Sigma^{0}_{1}$).
  • Cautious‑Scientist AI (bengio2024cautious): the basic trace is a finite Bayesian‑update ledger (decidable, so $\Delta^{0}_{1}$); if halting‑probability calls are additionally permitted, then the checker becomes one‑sided enumerable, hence $\Sigma^{0}_{1}$.
@@
2. Post’s theorem lifts the halting problem. post1944recursively showed that for any oracle $O\in\Sigma^{0}_{k}$ the OO‑relative halting set
@@
is not decidable by any total OO‑oracle Turing machine.
@@
[…]
@@
Appendix H Proof of Proposition 5